# Sensitivity and topology of exceptional rings in nonlinear non-Hermitian planar optical microcavities

### v1

@@ I Introduction
Non-Hermitian systems have attracted significant attention in recent years. In contrast to Hermitian systems they exhibit potentially complex eigenvalues and non-orthogonal eigenvectors enabling the investigation of striking and often counter-intuitive phenomena. Exceptional points are singularities at which two or more complex eigenvalues coalesce and their respective eigenvectors become parallel [1, 2, 3, 4, 5, 6]. The number of modes $n$ that coalesce at an exceptional point defines the order $n$ of the exceptional point. Exceptional points have been investigated in various platforms, including microwave resonators [7, 8, 9], atomic systems [10], plasmonic nanostructures [11, 12], optical waveguides [13], microresonators [14, 15, 16], and non-reciprocal systems [17]. The coalescence of eigenvectors often leads to counterintuitive behavior close to the exceptional point such as loss-induced transparency [18], loss-induced suppression and revival of lasing [19, 20], unidirectional invisibility, and reflectivity exceeding unity [13, 21]. Unlike diabolic points in Hermitian systems, the eigenvalue splitting scales with the $n^{\mathrm{th}}$-root of the perturbation strength close to an exceptional point of order $n$, rendering them highly sensitive [22, 23]. Although the discussion about noise enhancement remains ongoing, this sensitivity makes exceptional points promising candidates for new sensor designs [24, 25, 26]. Recent studies show that nonlinear systems provide excellent platforms to manipulate exceptional points and their characteristic features. Prominent examples include coupled optical and electrical resonators, and hybrid light-and-matter quasiparticles such as exciton-polaritons. Notable phenomena in these systems include mode-switching around exceptional points in bistable domains [27, 28], abrupt global changes in the stability due to coinciding exceptional points [29], phase-transitions [30, 31], robust wireless power transfer [32], as well as saturable gain and Kerr-nonlinearity induced shifting [33, 34] and rotation of exceptional points [34]. Remarkably, exceptional-like points can also be studied in nonlinear Hermitian systems [35].
@@
Figure 1: Concept of exceptional ring in planar microcavities with circular dichroism and spin-orbit coupling. (a) Sketch of planar resonator with circular dichroism in the form $\Gamma_{+}\neq\Gamma_{-}$ and finite TE-TM splitting. The complex-valued in-plane resonator dispersions, $\mu(\mathbf{k})$, host an ER centered at $k=0$ in reciprocal space. Nonlinearity splits that ring into mulitple rings (marked as black donuts). (b,d) Isotropic in-plane dispersions of the circularly polarized modes split into TE and TM modes by their coupling. (c,e) In the ER regime, when circular dichroism is introduced, $\Gamma_{+}\neq\Gamma_{-}$, TE-TM splitting is negated inside an exceptional ring (ER) forming in k-space at $|\mathbf{k}_{\mathrm{ER}}|=\sqrt{|\gamma|/|\Delta_{\mathrm{LT}}|}=1/\mathrm{\upmu m}$; the dispersions become circularly/elliptically polarized as shown by their $S_{3}$ Stokes vector element. Dispersions are rotationally invariant in $k$-space and $y$-axes in (b,c) are relative to the dispersion minimum.
@@
While the interplay of nonlinear and non-Hermitian physics is fascinating, in the nonlinear regime it is not always easy to tune the system to the vicinity of an isolated EP [34]. In contrast to distinct pairs of exceptional points, exceptional rings (ERs) form a continuous, gapless loop of coalescing eigenvectors and eigenvalues and offer a promising platform to drive a system to an exceptional point. It was shown that ERs may emerge from diabolic points in reciprocal space of non-Hermitian spin-orbit coupled systems [36]. To date, they have been studied in helical waveguide arrays [37], high-spin ultracold atoms [38], mechanical and thermal systems [39, 40, 41, 41], honeycomb [42] and superconducting Lieb lattices [43], photonic and sonic crystals [44, 45, 46, 47], as well as the excitation spectrum of light-matter systems [30]. In Ref. [48] a design for an ER-laser with pseudochiral optical response was presented in active plasmonic metasurfaces. By tweaking the gain and loss profiles, ERs can be tuned into different one-dimensional contours [49]. In Ref. [50] topological phase transitions between ERs with opposite topological charges were investigated. Recently, also the interplay of an ER and the nonlinear Hall effect was investigated [51].
@@ II Exceptional ring in planar resonators with spin-orbit coupling
We start with the linear problem where we find that our microcavity shows the interesting feature of a natural ER. Only two basic ingredients are required for this observation: (i) TE-TM splitting and (ii) different losses for left- and right-circularly polarized field components ($\Gamma_{+}\neq\Gamma_{-}$), a specific form of circular dichroism. We note that TE-TM splitting is typically intrinsically present (the specific functional form is not important here). Circular dichroism may be introduced, for example, through gyroscopic materials, assuming degeneracy of the resonances in the two circular components for the scalar case at zero TE-TM splitting [57, 58]. If non-degenerate, the modes would have to be tuned back into resonance for example by applying an external $B$-field [59, 60] or external off-resonant pumping in active semiconductor materials [61, 62]. As a specific realization, in much of the following, we study exciton polaritons in a semiconductor microcavity. The polariton amplitudes (or wave functions) of the lower polariton branch are denoted by $\psi$; they depend on the polariton wave vector k. The two circular polarization states (right and left circularly polarized) are indicated by the subscricts, $\psi_{+}$ and $\psi_{-}$. The physics of interest comes from the (in general different) decay rates in the two polarization channels and the coupling of the two channels.
We model the linear system by an effective $2\times 2$ Hamiltonian and solve the corresponding eigenvalue problem for each $\mathbf{k}$ on a discretized 2D $k$-space grid
$\begin{bmatrix}\frac{\hbar^{2}\textbf{k}^{2}}{2m}+i\Gamma_{+}&\Delta_{\mathrm{LT}}(ik_{x}+k_{y})^{2}\\ \Delta_{\mathrm{LT}}(ik_{x}-k_{y})^{2}&\frac{\hbar^{2}\textbf{k}^{2}}{2m}+i\Gamma_{-}\end{bmatrix}\begin{bmatrix}\psi_{+}\\ \psi_{-}\end{bmatrix}=\mu\begin{bmatrix}\psi_{+}\\ \psi_{-}\end{bmatrix}\,.$ (1)
@@
In the linear regime, the eigenvalue problem (1) can be solved straightforwardly. In Figure 1(b,d) the real and imaginary parts of the solutions in the linear isotropic regime (and $\gamma=0$) show the well-established TE and TM frequency splitting and equal losses along $k_{x}$. The displayed results are rotationally invariant in $k$-space. The coloring of the lines indicates the $S_{3}=\frac{|\psi_{+}|^{2}-|\psi_{-}|^{2}}{|\psi_{+}|^{2}+|\psi_{-}|^{2}}$ Stokes vector element for the respective eigenvector |$\ket{\psi}=\begin{bmatrix}\psi_{+}&\psi_{-}\end{bmatrix}^{\mathrm{T}}$ as a measure for the degree of circular polarization. Left and right circular, as well as linear polarizations are given by $S_{3}=-1$, $S_{3}=1$, and $S_{3}=0$, respectively.
In Figure 1(c,e) the respective solutions are shown in the linear non-Hermitian regime with circular dichroism ($\gamma=0.1~{}\mathrm{ps^{-1}}$). In this regime an ER with radius $k_{\mathrm{ER}}=|\mathbf{k}_{\mathrm{ER}}|=\sqrt{\nicefrac{{|\gamma|}}{{|\Delta_{\mathrm{LT}}|}}}=1~{}\mathrm{\upmu m^{-1}}$ is formed in the dispersion (the displayed results are rotationally symmetric in $k$-space). The mode splitting is canceled inside the ER where the two solutions become circularly polarized [cf. Fig. 1(c)]. The radius of the ER is determined by the ratio of TE-TM splitting $\Delta_{\mathrm{LT}}$ and degree of circular dichroism $\gamma$. Specific linewidths only matter for the visual appearance of the ring; the dispersion and spectral properties are best visible when the lines close to the ER radius do not overlap too much.
@@
In contrast to the eigenvalue problem in Eq. (1) this equation describes the field dynamics in two-dimensional real-space with $\psi_{\pm}=\psi_{\pm}(\mathbf{r},t)$, which is driven by an external coherent field $R_{\pm}=R_{\pm}(\mathbf{r},t)$. Consistent with the $k$-space representation in Eq. (1), the TE-TM splitting operator is defined as $J_{\pm}={\Delta_{\mathrm{LT}}}(\partial_{x}\mp i\partial_{y})^{2}$. To extract the system response that shows the dispersions, we pulse the system with a short resonant pulse $R_{\pm}(\textbf{r},t)=R_{0}\mathrm{exp}\left(-\frac{\textbf{r}^{2}}{p_{\mathrm{w}}^{2}}\right)\mathrm{exp}\left(-i\omega t\right)\mathrm{exp}\left(\frac{-(t-t_{0})^{2}}{\sigma^{2}}\right)$, separately in each polarization channel in two independent time evolution calculations. The calulations are performed with our CUDA compatible open-source solver for 2D nonlinear Schrödinger equations "PHOENIX" [67] and the code for the specific application shown here is included in the PHOENIX Github [68]. In Figure 2 we plot the diagonal elements of the response tensor, $\chi_{{}_{++}}=\frac{\mathcal{F}\left(\psi_{+}\right)}{\mathcal{F}\left(R_{+}\right)}$ and $\chi_{{}_{--}}=\frac{\mathcal{F}\left(\psi_{-}\right)}{\mathcal{F}\left(R_{-}\right)}$, which dominate the system response for small TE-TM splitting. Here, $\mathcal{F}(X)$ is the Fourier transform of the respective quantity $X$ into $k$- and $\omega$-space.
The resulting dispersions in Fig. 2 show perfect agreement with the results of the eigenvalue problem, Eq. (1). Both energy and linewidth behavior in the time-dependent calculations agree with the results of our reduced 2x2 model as displayed in Fig. 1(c,e). We note that, in principle, for suitably tailored $\gamma(\mathbf{k})$ or for non-$k$-isotropic systems, different shapes of the exceptional structure in 2d $k$-space can be realized, allowing for the formation of exceptional ellipses and more complex exceptional curves. In Appendix A, for the specific realization in an exciton polariton system, we explicitly show that off-resonant pumping preserves the assumed $k$-independence of the polarization dependent gain difference, $\gamma$, inferred in the present study.
@@ III Nonlinearity-induced third-order exceptional ring
with $M_{\pm}=\frac{\hbar^{2}\textbf{k}^{2}}{2m}+i\Gamma_{\pm}+g_{\mathrm{c}}|\psi_{\pm}|^{2}+g_{\mathrm{x}}|\psi_{\mp}|^{2}\pm\Delta$. Here, $g_{\mathrm{c}}$ and $g_{\mathrm{x}}$ define the Kerr-like nonlinearity and cross-mode interaction strength, respectively. We first focus on the degnerate case of $\Delta=0$ for which the fully analytic solutions of the nonlinear eigenvalue problem are of tractable length (see Appendix B). Later, through $\Delta\neq 0$ we will include finite mode detuning of different origins; examples we discuss are through magnetic field induced Zeeman splitting and electric field tunable Rashba-Dresselhaus splitting.
@@
To solve nonlinear eigenvalue problems, several approaches have been proposed, including the polar representation [74, 75] or Stokes parameters and momentum operators incorporation [76, 77, 78, 79]. However, these methods are often limited to specific Hamiltonian structures. More recently, iterative techniques based on self-consistent field approaches have been applied to non-Hermitian nonlinear eigenvalue problems [28, 80]. Despite their utility, iterative methods, such as the Newton-Raphson method show convergence issues in bistable regimes close to the exceptional point. In Ref. [52], we provide the general solutions (with analytical expressions in the case of zero detuning) of the eigenvalue problem (3); the solution approach used was inspired by the separability eigenvalue equations [81] as discussed in more detail in Appendix B. The resulting solutions apply to any real normalization $|\psi_{+}|^{2}+|\psi_{-}|^{2}=N^{2}$. The given solutions are valid also for the left eigenvalue problem under complex conjugation as explained in detail in Appendix B. Our analytical solutions are validated below using them as initial conditions for the Newton-Raphson method. Here, we scale the solutions to norm $N=1$, and introduce $\alpha=\nicefrac{{(g_{\mathrm{c}}-g_{\mathrm{x}})}}{{2}}$ as a measure for nonlinearity.
@@
Figure 3: Exceptional rings in the complex-valued energy spectra in nonlinear regime at zero detuning ($\Delta=0$). (a) Real and (b) imaginary parts of energy spectra of the nonlinear eigenvalue problem. The respective central cuts are depicted in (c-d). The energy spectrum hosts two exceptional rings, one with radius $|\textbf{k}|^{2}=\nicefrac{{|\gamma|}}{{|\Delta_{\mathrm{LT}}|}}=1~{}\mathrm{\upmu m}^{-1}$ (solid blue dot) and a second ring with radius $|\textbf{k}|^{2}=\nicefrac{{\sqrt{\gamma^{2}+\alpha^{2}}}}{{|\Delta_{\mathrm{LT}}|}}$ (solid violet dot). Between these two rings [marked in red in section B in panel (c-d)] four solutions are found. Inside the inner (outside the outer) ring two circularly (linearly) polarized solutions are found [section A and C in panel (c)]. The circles mark points where a solution crosses the boundary between two segments without tracing through an exceptional ring. The results displayed are rotational-symmetric in $k$-space.

### v2

@@ I Introduction
Non-Hermitian systems have attracted significant attention in recent years. In contrast to Hermitian systems they exhibit potentially complex eigenvalues and non-orthogonal eigenvectors enabling the investigation of striking and often counter-intuitive phenomena. Exceptional points are singularities at which two or more complex eigenvalues coalesce, and their respective eigenvectors become parallel [1, 2, 3, 4, 5, 6, 7]. The number of modes $n$ that coalesce at an exceptional point defines the order $n$ of the exceptional point. Exceptional points have been investigated in various platforms, including microwave resonators [8, 9, 10], atomic systems [11], plasmonic nanostructures [12, 13], optical waveguides [14], microresonators [15, 16, 17], and non-reciprocal systems [18]. The coalescence of eigenvectors often leads to counterintuitive behavior close to the exceptional point such as loss-induced transparency [19], loss-induced suppression and revival of lasing [20, 21], unidirectional invisibility, and reflectivity exceeding unity [14, 22]. Unlike diabolic points in Hermitian systems, the eigenvalue splitting scales with the $n^{\mathrm{th}}$-root of the perturbation strength close to an exceptional point of order $n$, rendering them highly sensitive [23, 24]. Although the discussion about noise enhancement remains ongoing, this sensitivity makes exceptional points promising candidates for new sensor designs [25, 26, 27]. Recent studies show that nonlinear systems provide excellent platforms to manipulate exceptional points and their characteristic features. Prominent examples include coupled optical and electrical resonators, and hybrid light-and-matter quasiparticles such as exciton-polaritons. Notable phenomena in these systems include mode-switching around exceptional points in bistable domains [28, 29], abrupt global changes in the stability due to coinciding exceptional points [30], phase-transitions [31, 32], robust wireless power transfer [33], as well as saturable gain and Kerr-nonlinearity induced shifting [34, 35] and rotation of exceptional points [35]. Remarkably, exceptional-like points can also be studied in nonlinear Hermitian systems [36].
@@
Figure 1: Concept of exceptional ring in planar microcavities with circular dichroism and spin-orbit coupling. (a) Sketch of planar resonator with circular dichroism in the form $\Gamma_{+}\neq\Gamma_{-}$ and finite TE-TM splitting. The complex-valued in-plane resonator dispersions, $\mu(\bm{k})$, host an ER centered at $k=0$ in reciprocal space. Nonlinearity splits that ring into mulitple rings (marked as black donuts). (b,d) Isotropic in-plane dispersions of the circularly polarized modes split into TE and TM modes by their coupling. (c,e) In the ER regime, when circular dichroism is introduced, $\Gamma_{+}\neq\Gamma_{-}$, TE-TM splitting is substracted inside an exceptional ring (ER) forming in $k$-space at $|\bm{k}_{\mathrm{ER}}|=\sqrt{|\gamma|/|\Delta_{\mathrm{LT}}|}=1/\mathrm{\upmu m}$; the dispersions become circularly/elliptically polarized as shown by their $S_{3}$ Stokes vector element. Dispersions are rotationally invariant in $k$-space and $y$-axes in (b,c) are relative to the dispersion minimum.
@@
While the interplay of nonlinear and non-Hermitian physics is fascinating, in the nonlinear regime it is not always easy to tune the system to the vicinity of an isolated EP [35]. In contrast to distinct pairs of exceptional points, exceptional rings (ERs) form a continuous, gapless loop of coalescing eigenvectors and eigenvalues and offer a promising platform to drive a system to an exceptional point. It was shown that ERs may emerge from diabolic points in reciprocal space of non-Hermitian spin-orbit coupled systems [37]. To date, they have been studied in helical waveguide arrays [38], high-spin ultracold atoms [39], mechanical and thermal systems [40, 41, 42, 42], honeycomb [43] and superconducting Lieb lattices [44], photonic and sonic crystals [45, 46, 47, 48, 49, 50, 51], as well as the excitation spectrum of light-matter systems [31]. In Ref. [52] a design for an ER-laser with pseudochiral optical response was presented in active plasmonic metasurfaces. By tweaking the gain and loss profiles, ERs can be tuned into different one-dimensional contours [53]. In Ref. [54] topological phase transitions between ERs with opposite topological charges were investigated. Recently, also the interplay of an ER and the nonlinear Hall effect was investigated [55].
@@ II Exceptional ring in planar resonators with spin-orbit coupling
We start with the linear problem where we find that our microcavity shows the interesting feature of a natural ER. Only two basic ingredients are required for this observation: (i) TE-TM splitting and (ii) different losses for left- and right-circularly polarized field components ($\Gamma_{+}\neq\Gamma_{-}$), a specific form of circular dichroism. We note that TE-TM splitting is typically intrinsically present (the specific functional form is not important here). Circular dichroism may be introduced, for example, through gyroscopic materials, assuming degeneracy of the resonances in the two circular components for the scalar case at zero TE-TM splitting [61, 62]. If non-degenerate, the modes would have to be tuned back into resonance for example by applying an external magnetic field [63, 64] or external off-resonant pumping in active semiconductor materials [65, 66]. As a specific realization, in much of the following, we study exciton polaritons in a semiconductor microcavity. The polariton amplitudes (or wave functions) of the lower polariton branch are denoted by $\psi$; they depend on the polariton wave vector $\bm{k}$. The two circular polarization states (right and left circularly polarized) are indicated by the subscricts, $\psi_{+}$ and $\psi_{-}$. The physics of interest comes from the (in general different) decay rates in the two polarization channels and the coupling of the two channels.
We model the linear system by an effective $2\times 2$ Hamiltonian and solve the corresponding eigenvalue problem for each $\bm{k}$ on a discretized 2D $k$-space grid
$\begin{bmatrix}\frac{\hbar^{2}\bm{k}^{2}}{2m}+i\Gamma_{+}&\Delta_{\mathrm{LT}}(ik_{x}+k_{y})^{2}\\ \Delta_{\mathrm{LT}}(ik_{x}-k_{y})^{2}&\frac{\hbar^{2}\bm{k}^{2}}{2m}+i\Gamma_{-}\end{bmatrix}\begin{bmatrix}\psi_{+}\\ \psi_{-}\end{bmatrix}=\mu\begin{bmatrix}\psi_{+}\\ \psi_{-}\end{bmatrix}\,$. (1)
Here, for sufficiently small $|\bm{k}|$, we assume parabolic in-plane dispersions for the left- and right-circularly polarized modes in absence of spin-orbit interaction ($\Delta_{\mathrm{LT}}=0$) and set the dispersion minimum to zero. The dispersion curvature is given by the effective mass $m=10^{-4}m_{\mathrm{e}}$, with $m_{\mathrm{e}}$ being the free electron mass. $\psi_{\pm}$ describes the field of the right- and left-circularly polarized modes with losses $\Gamma_{\pm}$, respectively. As a measure for the circular dichroism we define $\gamma=\nicefrac{{|(\Gamma_{+}-\Gamma_{-})|}}{{2}}$. The mode coupling, TE-TM splitting, is introduced by the off-diagonal elements, with the factor $(ik_{x}+k_{y})^{2}$ signifying a spin-orbit coupling with an exchange of $2\hbar$ between the orbital motion and the polaritonic spin (circular polarization) states. The coupling strength is given by $\Delta_{\mathrm{LT}}=0.1~\mathrm{meV\upmu m^{2}}$ and here assumed to be independent of the magnitude of the wavevector, $|\bm{k}|$. Such a dependence depends on the structural details of the system, and does not affect the occurance of an ER as long as the total TE-TM splitting is a monotonically increasing function of $|\bm{k}|$ at least up to the $|\bm{k}|$ at which the ER is observed [67, 68, 69, 70]. In other words, the specific functional form of the dispersion is not important for the observation of the ER and even in absence of $k$-space isotropy a deformed ER structure is found. Note that while here we focus on the lossy case of negative $\Gamma_{\pm}$ in the following, the eigenvalue problem also shows an ER for overall gain.
@@
In Figure 1(c,e) the respective solutions are shown in the linear non-Hermitian regime with circular dichroism ($\gamma=0.1~\mathrm{ps^{-1}}$) with $\{\Gamma_{+},\Gamma_{-}\}=\{-0.2,-0.4\}~\mathrm{ps^{-1}}$. In this regime an ER with radius $k_{\mathrm{ER}}=|\bm{k}_{\mathrm{ER}}|=\sqrt{\nicefrac{{|\gamma|}}{{|\Delta_{\mathrm{LT}}|}}}=1~\mathrm{\upmu m^{-1}}$ is formed in the dispersion (the displayed results are rotationally symmetric in $k$-space). The mode splitting is canceled inside the ER where the two solutions become circularly polarized [cf. Fig. 1(c)]. The radius of the ER is determined by the ratio of TE-TM splitting $\Delta_{\mathrm{LT}}$ and degree of circular dichroism $\gamma$. Specific linewidths only matter for the visual appearance of the ring; the dispersion and spectral properties are best visible when the lines close to the ER radius do not overlap too much.
@@
In contrast to the eigenvalue problem in Eq. (1) this equation describes the field dynamics in two-dimensional real-space with $\psi_{\pm}=\psi_{\pm}(\mathbf{r},t)$, which is driven by an external coherent field $R_{\pm}=R_{\pm}(\mathbf{r},t)$. Consistent with the $k$-space representation in Eq. (1), the TE-TM splitting operator is defined as $J_{\pm}={\Delta_{\mathrm{LT}}}(\partial_{x}\mp i\partial_{y})^{2}$. To extract the system response that shows the dispersions, we pulse the system with a short resonant pulse $R_{\pm}(\textbf{r},t)=R_{0}\mathrm{exp}\left(-\frac{\textbf{r}^{2}}{p_{\mathrm{w}}^{2}}\right)\mathrm{exp}\left(-i\omega t\right)\mathrm{exp}\left(\frac{-(t-t_{0})^{2}}{\sigma^{2}}\right)$, separately in each polarization channel in two independent time evolution calculations. The calulations are performed with our CUDA compatible open-source solver for 2D nonlinear Schrödinger equations "PHOENIX" [71] and the code for the specific application shown here is included in the PHOENIX Github [72]. In Fig. 2 we plot $\mathrm{Im}\left(\chi\right)$ for the diagonal elements of the response tensor, $\chi_{{}_{++}}=\frac{\mathcal{F}\left(\psi_{+}\right)}{\mathcal{F}\left(R_{+}\right)}$ and $\chi_{{}_{--}}=\frac{\mathcal{F}\left(\psi_{-}\right)}{\mathcal{F}\left(R_{-}\right)}$, which dominate the system response for small TE-TM splitting. Here, $\mathcal{F}(X)$ is the Fourier transform of the respective quantity $X$ into $k$- and $\omega$-space.
The resulting dispersions in Fig. 2 show perfect agreement with the results of the eigenvalue problem, Eq. (1). Both energy and linewidth behavior in the time-dependent calculations agree with the results of our reduced 2x2 model as displayed in Fig. 1(c,e). We note that, in principle, for suitably tailored $\gamma(\bm{k})$ or for non-$k$-isotropic systems, different shapes of the exceptional structure in 2D $k$-space can be realized, allowing for the formation of exceptional ellipses and more complex exceptional curves. In Appendix A, for the specific realization in an exciton polariton system, we explicitly show that off-resonant pumping preserves the assumed $k$-independence of the polarization dependent gain difference, $\gamma$, inferred in the present study.
@@ III Nonlinearity-induced third-order exceptional ring
with $M_{\pm}=\frac{\hbar^{2}\bm{k}^{2}}{2m}+i\Gamma_{\pm}+g_{\mathrm{c}}|\psi_{\pm}|^{2}+g_{\mathrm{x}}|\psi_{\mp}|^{2}\pm\Delta$. Here, $g_{\mathrm{c}}$ and $g_{\mathrm{x}}$ define the Kerr-like nonlinearity and cross-mode interaction strength, respectively. We first focus on the degnerate case of $\Delta=0$ for which the fully analytic solutions of the nonlinear eigenvalue problem are of tractable length (see Appendix B). Later, through $\Delta\neq 0$ we will include finite mode detuning of different origins; examples we discuss are through magnetic field induced Zeeman splitting and electric field tunable Rashba-Dresselhaus splitting.
@@
To solve nonlinear eigenvalue problems, several approaches have been proposed, including the polar representation [78, 79] or Stokes parameters and momentum operators incorporation [80, 81, 82, 83]. However, these methods are often limited to specific Hamiltonian structures. More recently, iterative techniques based on self-consistent field approaches have been applied to non-Hermitian nonlinear eigenvalue problems [29, 84]. Despite their utility, iterative methods, such as the Newton-Raphson method show convergence issues in bistable regimes close to the exceptional point. In Ref. [56], we provide the general solutions (with analytical expressions in the case of zero detuning) of the eigenvalue problem (3); the solution approach used was inspired by the separability eigenvalue equations [85] as discussed in more detail in Appendices B and C. The resulting solutions apply to any real normalization $|\psi_{+}|^{2}+|\psi_{-}|^{2}=N^{2}$. The given solutions are valid also for the left eigenvalue problem under complex conjugation (Appendix B). Our analytical solutions are validated using them as initial conditions for the Newton-Raphson method. Here, we scale the solutions to norm $N=1$, and introduce $\alpha=\nicefrac{{(g_{\mathrm{c}}-g_{\mathrm{x}})}}{{2}}$ as a measure of nonlinearity.
@@
Figure 3: Exceptional rings in the complex-valued energy spectra in nonlinear regime at zero detuning ($\Delta=0$). Central cuts of the (a) real and (b) imaginary parts of energy spectra of the linear eigenvalue problem (neglecting the $k^{2}$ dispersion). The respective surfaces are depicted in the insets with the exceptional ring (ER) marked in black. Central cuts of the (c) real and (d) imaginary parts of energy spectra of the nonlinear eigenvalue problem. The respective surfaces are depicted in (e-f). In contrast to the linear regime, the energy spectrum in the nonlinear regime hosts two exceptional rings, one with radius $|\bm{k}|^{2}=\nicefrac{{|\gamma|}}{{|\Delta_{\mathrm{LT}}|}}=1~\mathrm{\upmu m}^{-1}$ (solid blue dot and labeled by EAR) and a second ring with radius $|\bm{k}|^{2}=\nicefrac{{\sqrt{\gamma^{2}+\alpha^{2}}}}{{|\Delta_{\mathrm{LT}}|}}$ (solid violet dot and labeled by EXR). Between these two rings [marked in red in section B in panel (c-d)] four solutions are found. Inside the inner (outside the outer) ring two circularly (linearly) polarized solutions are found [section A and C in panel (c,d)]. The circles mark points where a solution crosses the boundary between two segments without tracing through an exceptional ring and the cross marks $k=0$. These markers are used in Fig 4(b) to analyze the bifurcation behavior near the two rings. The results displayed are rotationally symmetric in $k$-space. Simulation parameters: $\{\alpha,g_{\mathrm{c}},g_{\mathrm{x}}\}=\{0.125,0.25,0\}~\mathrm{meV\upmu m}^{2}$.